If v(t) = 10 cos(120πt+30o
) and i(t) = 6 cos(120πt+60o
), determine the average power, rms value of v(t),
and the power factor
1
Expert's answer
2020-11-13T10:51:20-0500

Given "v(t) = 10\cos(120\pi t+30\degree)"

Given "i(t) =6\cos(120\pi t+60)"


"Vrms=\frac{10}{\sqrt2}=7.071V" (rms value of "v(t)."

"Irms = \frac{6}{\sqrt2}= 4.243A"


Phase difference between current and voltage "(\phi)=60-30=30\degree lead"


Powerfactor"=\cos\phi=\cos30=0.866 \space lead"


Average power"=Vrms\times Irms\times\cos\phi=7.071\times 4.243\times0.866= 25.98\space W"